in the equation
x^2-x+2x=6, X=Y-1. what is one possible value of Y?
A. 2
B. 3
C. -3
D. 0

Answer by Boreal(15235)

Substitute or plug in answers
x^2+x-6=0, combining x s
(x+3)(x-2)=0
x=-3 or 2
y is x+1, so y = -2 or 3
B.